# Coward: Collision-based Watermark for Proactive Federated Backdoor Detection

### v3

@@ 5.2 Main Results
CIFAR-10 Vanilla 5.5 32.7 97.0 91.7 53.5 44.4 92.8 91.9 3.0 24.2 96.6 91.9 77.9 64.0 30.2 91.8 30.5 29.9 96.3 91.7 93.0 34.7 4.9 91.9 100.0 3.3 4.8 91.5
PGD 5.5 32.7 97.3 91.7 53.5 44.6 92.7 91.9 2.0 24.5 96.4 91.8 75.5 61.0 32.3 91.9 12.0 31.8 97.8 91.7 92.0 34.1 5.5 91.6 97.5 4.2 5.9 91.4
Neurotoxin 6.0 32.7 96.4 91.7 15.5 43.1 96.0 91.8 2.5 24.3 95.5 91.8 69.2 58.5 41.1 91.9 30.0 29.9 95.5 91.4 52.5 31.7 79.1 91.4 99.5 7.8 4.1 91.6
Chameleon 8.0 32.4 94.1 91.8 46.5 44.1 88.4 91.7 3.5 23.6 93.4 91.8 39.0 36.5 62.4 91.3 10.0 32.2 94.9 91.6 87.0 33.8 8.4 91.3 94.5 9.1 26.1 91.4
CIFAR-100 Vanilla 6.5 32.6 92.9 69.5 0.0 0.0 88.9 69.6 3.5 23.7 92.8 69.6 48.7 44.4 53.1 69.4 7.5 32.5 93.0 69.4 91.0 45.3 9.3 68.9 100.0 1.9 0.8 68.8
PGD 6.5 32.6 92.9 69.6 0.0 0.0 92.0 69.7 2.5 23.4 92.7 69.7 46.7 43.0 49.3 69.6 2.0 33.1 93.4 69.4 88.0 45.7 16.1 68.9 100.0 5.8 0.8 69.0
Neurotoxin 7.0 32.6 91.4 69.4 0.0 0.0 91.7 69.4 4.5 23.2 92.7 69.7 51.3 42.7 47.4 69.6 2.5 33.1 92.9 69.3 96.0 46.9 2.3 69.2 100.0 9.4 0.8 69.0
Chameleon 0.0 33.3 85.6 69.3 0.0 0.0 90.7 69.5 5.0 24.3 86.3 69.6 41.0 34.7 60.7 69.4 0.0 33.3 87.1 69.4 98.5 47.5 16.0 66.7 98.2 7.3 0.9 69.1
EMNIST Vanilla 8.5 32.4 100.0 99.7 12.0 35.7 100.0 99.7 1.0 24.6 100.0 99.7 24.0 37.3 100.0 99.7 44.5 28.3 100.0 99.7 88.3 45.7 21.4 99.7 100.0 3.6 0.0 99.7
PGD 10.0 32.2 99.9 99.7 11.0 35.2 100.0 99.7 2.0 11.7 100.0 99.7 44.0 46.9 99.8 99.7 37.0 31.4 33.0 99.7 88.5 46.0 25.3 99.6 100.0 4.3 0.0 99.7
Neurotoxin 10.5 31.9 99.9 99.7 2.5 36.3 100.0 99.7 1.5 12.0 100.0 99.7 38.5 43.1 100.0 99.7 44.5 17.2 96.8 99.7 88.6 45.3 16.0 98.9 99.0 3.6 0.1 99.7
Chameleon 12.5 31.9 100.0 99.7 21.5 37.2 100.0 99.7 3.0 24.8 100.0 99.6 37.5 42.7 100.0 99.7 98.0 22.3 29.3 99.7 91.5 51.7 39.2 99.6 99.5 2.4 15.4 99.7
Resistance to FL Non-i.i.d. We evaluate all baseline methods under varying degrees of non-i.i.d. data across three datasets, using a vanilla backdoor setting with a fixed spot-pattern trigger and a learning rate matching that of benign clients. As shown in Table II, Coward consistently achieves the best overall performance across all levels of heterogeneity, with TPRs above 95% and FPRs below 10.5% in all cases. Besides the overall conclusion, we draw several key observations: (1) Coward significantly reduces FPR across all settings compared to BackdoorIndicator, highlighting its advantage in mitigating OOD bias. A more detailed analysis is provided in Section 5.5. (2) Proactive methods significantly outperform passive ones under severe data heterogeneity. Taking the most challenging non-i.i.d. setting ($\alpha=0.3$) as an example, aside from our method, BackdoorIndicator also outperforms all passive baselines, achieving TPRs of 93% on CIFAR-10 and 96% on EMNIST. This result supports the robustness of proactive detection in heterogeneous environments. In contrast, passive methods exhibit significant performance degradation. For instance, the state-of-the-art passive method Flame sees its TPR drop from above 95% to below 10% on both CIFAR-100 and EMNIST, which aligns with our earlier analysis. (3) Some passive methods fail due to their unrealistic assumptions. For example, Rflbat performs poorly across all heterogeneity levels, likely because it relies only on the top two principal components of model parameters, losing critical information, especially under our challenging FL setup. Similarly, FoolsGold collapses on CIFAR-100, always judging all clients as benign (resulting in 0% TPR and 0% FPR). This failure stems from two factors: the high-dimensional 100-class prediction vector and its reliance on multi-attacker scenarios. These conditions lead to extremely low cosine similarity values, causing an all-as-benign judgment which is unhelpful.
TABLE IV: Detection performance under multi-client backdoor attack on CIFAR-10 dataset. Result highlighting (in boldface and in red) follows the same rules in Table II.
Strategy Atk_ratio MultiKrum FoolsGold Rflbat DeepSight Flame Indicator Coward (Ours)
@@
Uniform 30% 100.0 71.4 10.0 91.6 100.0 34.2 9.7 91.8 39.4 22.5 88.7 91.6 3.7 14.6 97.8 91.6 100.0 14.3 9.7 91.9 99.2 29.5 11.1 91.5 100.0 6.7 9.9 91.5
50% 3.0 97.0 98.4 90.8 100.0 27.1 10.1 91.8 31.3 0.0 91.9 91.4 1.1 24.0 98.2 91.4 21.5 57.3 99.2 91.4 99.1 16.3 12.2 91.2 100.0 8.2 9.9 91.2
70% 29.0 98.9 98.2 90.8 100.0 12.8 9.7 91.8 14.3 11.8 93.5 90.0 2.4 54.8 98.8 91.5 31.3 57.8 99.0 91.5 99.4 44.6 40.5 88.8 100.0 11.2 10.5 89.6
DBA 30% 100.0 57.4 9.8 91.6 100.0 32.9 9.5 91.9 28.0 15.4 81.9 91.5 5.6 12.4 90.6 91.5 100.0 14.3 9.5 91.9 98.2 18.4 10.1 91.8 100.0 9.7 9.7 91.3
50% 22.8 77.2 96.7 91.1 100.0 26.7 9.9 91.9 29.2 0.0 90.7 91.3 2.9 20.1 94.7 91.6 60.5 18.9 81.9 91.7 98.6 16.0 10.6 91.5 100.0 9.9 10.0 91.2
70% 31.1 94.1 97.7 90.1 100.0 12.8 9.6 91.8 9.5 2.8 95.3 88.8 3.4 44.7 96.0 90.9 42.0 34.8 96.5 91.3 98.8 15.8 11.9 89.4 100.0 11.2 10.3 89.6
NBA 30% 100.0 57.1 10.1 91.7 37.5 48.4 88.6 91.7 78.5 70.8 65.7 91.7 2.8 31.2 90.1 91.8 55.2 18.9 86.5 92.0 97.8 39.4 11.0 91.2 98.7 7.1 10.1 91.5
50% 63.7 36.3 91.5 91.6 33.3 54.6 93.4 91.6 16.5 15.8 85.2 91.4 1.4 48.1 93.9 91.7 19.2 40.6 94.8 91.1 93.7 38.2 32.6 91.5 98.5 15.0 10.1 91.5
70% 36.9 80.7 96.5 90.4 23.4 62.7 96.1 90.8 0.1 0.0 92.1 89.4 1.3 78.8 97.1 90.6 10.6 74.8 96.4 90.6 81.3 14.0 66.7 90.8 86.2 14.6 24.7 91.6
Resistance to Advanced Attacks. We evaluate all baselines under more challenging client-side attacks. Specifically, across all attack methods, the malicious client adopts a more imperceptible noise-blended trigger and reduced local learning rate. These advanced configurations significantly increase the stealthiness of the attack. As shown in Table III, our method achieves the best overall performance, with proactive detection showing strong robustness, while passive methods suffer substantial degradation. This highlights the advantage of the proactive paradigm in handling sophisticated attacks. (1) Among proactive methods, Coward consistently achieves the best performance, maintaining close to 100% TPR and sub-10% FPR across all advanced attack types, including state-of-the-art strategies such as Neurotoxin and Chameleon. BackdoorIndicator also performs strongly, with TPRs consistently exceeding 85% in most settings; however, it suffers from relatively high FPRs, reflecting its vulnerability to distributional bias. (2) For passive methods, performance degradation is anticipated under such stealthy attacks, since the attacker introduces only subtle changes to the model, which reduces its suspiciousness. As an instance, MultiKrum, which selects representative updates based on Euclidean distance, becomes highly ineffective, with TPRs falling below 12.5% across all stealthy attack settings. Flame, which filters out updates with large angular deviations, also suffers notable degradation, with TPRs dropping below 30.5% on CIFAR-10 dataset. In contrast, DeepSight shows comparatively better resilience, which perform consistently with TPRs over 40% in most advanced attack cases. This improvement most probably stems from its joint use of both parameter and prediction discrepancies.
Robustness to Multiple Attackers. Table IV presents results under multi-attacker scenarios. As shown in this table, proactive methods exhibit better overall performance across all three settings. This is expected as the detection mechanism of proactive approaches is independent of the number of attackers. Among them, our method achieves the best overall performance, consistently maintaining TPRs above 99% and FPRs below 15%. BackdoorIndicator also performs well but becomes less effective when the attacker proportion reaches 70%. In such cases, a temporary detection failure allows the increased number of attackers to rapidly reinforce the backdoor in the global model, resulting in a higher ASR. Moreover, the retained backdoor could be further maintained by the indicator’s own planting process due to the maintenance effect. Among passive baselines, FoolsGold performs best under Uniform and DBA attacks, achieving a TPR of 100%, benefiting from its assumption of highly similar, coordinated malicious updates. However, it fails in uncoordinated scenarios, where attackers pursue different targets. In such cases, the TPR drops below 40% as the method’s core assumption of shared attack objectives no longer holds. Outlier-based methods such as MultiKrum and Flame become ineffective when attackers control 70% of clients, with TPRs dropping under 42% in all cases. These methods assume malicious updates are outliers, which fails when adversaries form the majority among all clients.
TABLE V: Performance on different OOD planting sets. Result highlighting (in boldface and in red) follows the same rules in Table II.
Main Task Planting Set Indicator Coward
@@
CIFAR-10 EMNIST 98.0 24.0 10.2 91.6 99.5 7.3 9.7 91.8
CIFAR-100 97.0 14.2 10.0 91.7 99.5 6.7 9.9 91.2
NOISE 100.0 57.4 9.8 90.7 100.0 12.7 10.1 90.6
CIFAR-100 EMNIST 98.0 34.0 1.3 68.2 100.0 5.3 1.0 68.9
CIFAR-10 99.0 44.6 1.8 67.2 100.0 6.3 1.2 67.5
NOISE 98.0 60.0 1.5 67.9 100.0 4.4 1.0 67.8
EMNIST CIFAR-10 95.0 57.8 10.0 99.6 99.5 2.2 10.0 99.8
CIFAR-100 97.0 58.7 10.0 99.6 99.0 1.5 10.1 99.7
NOISE 100.0 73.5 10.0 99.6 99.0 4.8 10.2 99.6
5.3 Ablation Study
We investigate key factors that influence watermark planting and detection: (1) the choice of OOD planting set, (2) the pairing of watermark and backdoor triggers, and (3) the selection of detection thresholds. For a more comprehensive analysis, please refer to Appendices B and C.
Robustness to Choices of OOD Datasets. Table V presents detection performance across diverse OOD planting sets, including task-similar data, distinctly different data, and an extreme case with synthetic Gaussian noise. We observe that Coward consistently maintains low FPR across all settings, while Indicator’s performance degrades as the OOD dataset diverges from the main task. Using the most challenging main task on CIFAR-100 dataset as example, Coward’s FPR remains below 6.3% across all planting sets. In contrast, Indicator’s FPR increases sharply from 44.6% on CIFAR-10 to 60% with Gaussian noise. This confirms our earlier analysis of Indicator’s increased vulnerability to OOD bias. Overall, these results demonstrate that Coward offers improved robustness to the choice of OOD data, enabling more flexible deployment in real-world scenarios.
Robustness to Choices of Triggers. Table VI summarizes detection performance under various combinations of trigger types. The attacker’s trigger varies in increasing stealthiness: from a visible pixel pattern, to an invisible blend pattern, and finally to an implicit semantic pattern naturally selected from the dataset. The watermark trigger also spans a range of types, including four visible patterns (diagonal, square, triangular, and noise), as well as two invisible variants: a mosaic pattern and the sample-specific WaNet pattern. We see that Coward performs consistently well across all settings, achieving a TPR above 96.5% and an FPR below 17.3%. These results demonstrate our method’s strong robustness to different triggers and the generalizability of the multi-backdoor collision effect.
TABLE VI: Detection performance with different combinations of watermark trigger (S) and attackers’ client-side malicious trigger (A) on CIFAR-10.
[…]
Diagonal 100.0 6.1 9.8 91.6 99.5 6.6 5.2 91.7 98.5 17.3 0.0 91.8
Square 100.0 9.4 9.6 91.4 99.5 9.1 4.8 91.4 98.5 16.2 0.0 91.4
Triangle 99.5 4.6 9.8 91.6 98.5 6.2 7.8 91.7 99.0 12.3 0.0 90.8
Noise 99.5 6.1 9.9 91.5 99.0 6.7 7.9 91.5 99.5 7.3 0.0 91.4
Mosaic 96.5 10.6 18.0 91.6 99.5 12.0 4.6 91.3 99.0 14.2 0.0 90.7
WaNet 99.5 7.3 9.8 91.4 99.5 6.8 5.4 91.5 98.0 11.9 0.0 90.8
TABLE VII: Detection performance of Coward under different watermark-retention thresholds on the CIFAR-10 dataset.
[…]
1% 98.0 1.3 10.2 91.5
5% 100.0 2.6 10.1 91.5
10% 100.0 8.6 10.1 91.4
15% 100.0 12.0 9.9 91.5
20% 100.0 13.9 10.0 91.2
Robustness to Choices of Detection Thresholds. We further evaluate the impact of the detection threshold on the effectiveness of Coward. Since both attackers and benign clients may disrupt the watermark to some extent, the optimal threshold tends to fall within a relatively low range. Thus, we conduct threshold analysis within a mild range from 1% to 20%. As shown in Table VII, varying the detection threshold has minimal impact on TPR, which consistently remains above 98%. However, it progressively increases the FPR, rising from 1.0% to 13.9%. This increase is primarily due to client-side data heterogeneity. Certain benign clients, particularly those with severely skewed local distributions, may forget the watermark and exhibit significantly lower watermark accuracy than the majority. In contrast, malicious clients consistently show very low watermark accuracy. Empirically, we find that setting the threshold at 5% offers a reliable and practical guideline for a balanced TPR and FPR across diverse scenarios.
5.4 Resistance to Potential Adaptive Attack
TABLE VIII: Performance of Coward under the adaptive attack (AA). The subscripts $l$ and $g$ denote the ASR of the attacker’s backdoor evaluated on the malicious local model and the aggregated global model, respectively.
[…]
w./o. defense / / 91.2 95.6
w./o. AA. 100.0 5.0 10.1 98.9
AA. ($T=5$) 99.5 4.4 10.1 90.4
AA. ($T=1$) 62.5 5.4 22.9 73.9
@@ 5.4 Resistance to Potential Adaptive Attack
Potential Adaptive Attack Strategy. We consider adaptive attackers attempt to bypass our collision-based detection by injecting a similar watermark to preserve high watermark accuracy. Since the watermark target is chosen from a finite, known set, the attacker has a non-trivial chance of guessing it. To model this, we design a periodic guessing attack with switching period $T$, where the attacker updates the guessed label every $T$ rounds. Smaller $T$ increases the chance of early success but may cause unstable feedback, while larger $T$ offers more stable signals at the cost of later success.
@@
Results. As shown in Table VIII, the guessing attack appears to reduce the TPR but does not truly bypass the detection. For instance, when $T=1$, the TPR drops from 100% to 62.5%, indicating some disruption due to a correct guess. However, the attacker’s main objective is to increase the ASR on the aggregated global model. Notably, the global ASR remains low across all adaptive settings, reaching only 22.9% at most when $T=1$, which falls short to be successful. Upon further analysis, we find that the guessing attack inherently suffers from a fundamental contradiction, which we refer to as the local collision contradiction. Specifically, when the attacker injects an additional watermark with a different target label, it tends to induce a multi-backdoor conflict within the local model. This interference weakens the original backdoor and reduces the overall effectiveness of the attack. Overall, Coward remains resilient, as attackers struggle to maintain high ASR without compromising their own attack objectives.
@@ 5.5 Why is Our Method Effective?
Overall OOD Bias Analysis. We evaluate the overall bias level under vanilla FL, BackdoorIndicator, and our method, and further investigate its correlation with detection false positive rate. Specifically, the prediction bias is quantified by the standard deviation of class-wise prediction probabilities, since uniform predictions are ideal for OOD samples, and any bias toward a specific class results in non-uniformity, thereby increasing the variance. By aggregating the bias across all participating clients over all training epochs, we obtain the overall bias level throughout the training process. As results shown in Figure 7, it is evident that: (1) Our method yields the lowest bias distribution, validating the effectiveness of regulating the OOD mapping and highlighting the adverse effect of BackdoorIndicator’s random mapping strategy. (2) Even at comparable bias levels, our method achieves a lower FPR, demonstrating its robustness against OOD bias.
@@
Figure 7: A Holistic Quantification of OOD Bias and Its Impact on FPR. The left panel illustrates the OOD bias distribution, while the right panel shows the relationship between bias severity and the number of misjudged benign clients. Our method significantly reduces the OOD bias level, leading to a lower false positive rate, whereas the BackdoorIndicator increases the bias level and exhibits a high false positive frequency in highly biased regions.
Case Study. To intuitively explain how our method mitigates OOD bias, we present a set of representative detection cases of our method in Figure 8, ranging from an ideal non-bias scenario to increasingly biased conditions. Each case visualizes the prediction confusion matrices of both raw and watermarked OOD samples, revealing how local training and OOD bias impact the two injected mappings. Specifically, the top two rows show the prediction results of raw OOD samples, while the bottom two rows show those of watermarked samples. Each case includes a small subfigure illustrating the cumulative distribution across target classes. To aid intuitive understanding, we highlight that the attack target is class 0, and the watermark target is class 8. We elaborate on the case study by explaining both the ideal unbiased scenario and the biased scenario.:
  • Ideal Cases: Case 1 illustrates a malicious client (C0), where both the OOD and watermark mappings collapse into the attack target class 0, indicating severe OOD mapping distortion and a strong collision effect caused by the backdoor. This results in low watermark accuracy, enabling effective detection. In contrast, Case 2 represents an ideal benign client unaffected by OOD bias, where both the OOD (diagonal) and watermark mappings are fully preserved. Empirically, OOD bias seldom overrides backdoors but often impacts benign clients; Thus we examine biased cases for benign clients below.
@@
  • Biased Cases: Case 3 presents a mild OOD bias where the OOD mapping is largely retained but partially biased toward class 9, yet the watermark prediction remains dominated by class 8. This highlights the strength of the watermark, which consistently guides triggered OOD samples to the target class, even when raw samples are misclassified. In Case 4, under a stronger bias where OOD predictions collapse into two dominant classes, the watermark still holds the highest confidence for class 8, indicating resilience even under severe distributional shifts. Intuitively, after watermark planting, the model forms a direct association between the trigger pattern and the target label, making it less influenced by the background image and thus more robust to OOD bias. Case 5 presents an extreme scenario where the OOD mapping collapses to a single class that coincidentally aligns with the watermark target. These cases collectively indicate our robustness against different degrees of OOD bias.

### v4

@@ 5.3 Ablation Study
Table 6: Performance of Coward under different watermark-retention thresholds on the CIFAR-10 dataset.
@@
Robustness to Choices of Detection Thresholds. We further evaluate the impact of the detection threshold on the effectiveness of Coward. Since both attackers and benign clients may disrupt the watermark to some extent, the optimal threshold tends to fall within a relatively low range. Thus, we conduct threshold analysis within a mild range from 1% to 20%. As shown in Table 6, varying the detection threshold has minimal impact on TPR, which consistently remains above 98%. However, it progressively increases the FPR, rising from 1.0% to 13.9%. This increase is primarily due to client-side data heterogeneity. Certain benign clients, particularly those with severely skewed local distributions, may forget the watermark and exhibit significantly lower watermark accuracy than the majority. In contrast, malicious clients consistently show very low watermark accuracy. Empirically, we find that setting the threshold at 5% offers a reliable and practical guideline for a balanced TPR and FPR across diverse scenarios.
5.4 Resistance to Potential Adaptive Attack
Table 7: Performance of Coward under the adaptive attack (AA). The subscripts $l$ and $g$ denote the ASR of the attacker’s backdoor evaluated on the malicious local model and the aggregated global model, respectively.
@@ 5.4 Resistance to Potential Adaptive Attack
Potential Adaptive Attack Strategy. We consider adaptive attackers who attempt to bypass our collision-based detection by injecting a similar watermark to preserve high watermark accuracy. Since the watermark target is chosen from a finite, known set, the attacker has a non-trivial chance of guessing it. To model this, we design a periodic guessing attack with switching period $T$, where the attacker updates the guessed label every $T$ rounds. Smaller $T$ increases the chance of early success but may cause unstable feedback, while larger $T$ offers more stable signals at the cost of later success.
@@
Results. As shown in Table 7, the guessing attack appears to reduce the TPR but does not truly bypass the detection. For instance, when $T=1$, the TPR drops from 100% to 62.5%, indicating some disruption due to a correct guess. However, the attacker’s main objective is to increase the ASR on the aggregated global model. Notably, the global ASR remains low across all adaptive settings, reaching only 22.9% at most when $T=1$, which falls short of being successful. Upon further analysis, we find that the guessing attack inherently suffers from a fundamental contradiction, which we refer to as the local collision contradiction. Specifically, when the attacker injects an additional watermark with a different target label, it tends to induce a multi-backdoor conflict within the local model. This interference weakens the original backdoor and reduces the overall effectiveness of the attack. Overall, Coward remains resilient, as attackers struggle to maintain high ASR without compromising their own attack objectives.
@@ 5.5 Why is Our Method Effective?
Overall OOD Bias Analysis. We evaluate overall bias under vanilla FL, BackdoorIndicator, and our method, and further examine its correlation with the detection false positive rate. Specifically, the prediction bias is quantified by the standard deviation of class-wise prediction probabilities, since uniform predictions are ideal for OOD samples, and any bias toward a specific class results in non-uniformity, thereby increasing the variance. By aggregating the bias across all participating clients over all training epochs, we obtain the overall bias level. As the results shown in Figure 7 indicate, it is evident that: (1) Our method yields the lowest bias distribution, validating the effectiveness of regulating the OOD mapping and highlighting the adverse effect of BackdoorIndicator’s random mapping strategy. (2) Even at comparable bias levels, our method achieves a lower FPR, demonstrating its robustness against OOD bias.
@@
Figure 7: A Holistic Quantification of OOD Bias and Its Impact on FPR. The left panel shows the bias distribution, while the right panel shows the relationship between bias severity and the number of misjudged benign clients. Our method significantly reduces the OOD bias level, leading to fewer misjudgments, whereas the Indicator increases the bias level and exhibits a high false positive frequency in biased regions.
@@
Case Study. To intuitively explain how our method mitigates OOD bias, we present a set of representative detection cases of our method in Figure 8, ranging from an ideal unbiased scenario to increasingly biased conditions. Each case visualizes the prediction confusion matrices of both raw and watermarked OOD samples, revealing how OOD bias impact the two injected mappings. Specifically, the top two rows show the prediction results of raw OOD samples, while the bottom rows show those of watermarked samples. Each case includes a small subfigure illustrating the cumulative distribution across target classes. To aid intuitive understanding, we highlight that the attack target is class 0, and the watermark target is class 8. We elaborate on the case study by explaining both the ideal unbiased scenario and the biased scenario. (1) Ideal Cases: Case 1 illustrates a malicious client (C0), where both the OOD and watermark mappings collapse into the attack target class 0, indicating severe OOD mapping distortion and a strong collision effect caused by the backdoor. This results in low watermark accuracy, enabling effective detection. In contrast, Case 2 represents an ideal benign client unaffected by OOD bias, where both the OOD (diagonal) and watermark mappings are fully preserved. Empirically, OOD bias seldom overrides backdoors but often impacts benign clients. (2) Biased Cases: Case 3 presents a mild bias where the OOD mapping is largely retained but partially biased toward class 9, yet the watermark prediction remains dominated by class 8. This highlights the strength of the watermark, which consistently guides triggered samples to the target class, even when raw samples are misclassified. In Case 4, under a stronger bias where OOD predictions collapse into two dominant classes, the watermark still holds the highest confidence for class 8, indicating resilience even under severe distributional shifts. Intuitively, after watermark planting, the model forms a direct association between the trigger pattern and the target label, making it less influenced by the background image and thus more robust to OOD bias. Case 5 presents an extreme scenario where the OOD mapping collapses to a single class that coincidentally aligns with the watermark target. These cases collectively indicate the robustness of our method against different degrees of OOD bias.